# Leveraging multi-source and heterogeneous signals for fatigue detection

### v1

@@ Abstract
keywords:
@@
journal: Knowledge-Based Systems
[…]
[a]organization=Rowan University, addressline=201 Mullica Hill Rd, city=Glassboro, postcode=08028, state=NJ, country=USA
[…]
[b]organization=College of Mathematics and Statistics, Chongqing Jiaotong University, addressline=66 Xuefu Blvd, Nan’An, city=Chongqing, postcode=400074, country=China
{graphicalabstract}[Uncaptioned image]
{highlights}
We propose a practical, multi-source fatigue detection framework that addresses the challenge of leveraging high-fidelity knowledge from differently instrumented sensor domains to improve fatigue predication in real-world settings. The framework dynamically aligns heterogeneous signals across overlapping fatigue-related feature spaces, infers missing sensor modalities through targeted imputation, and continuously refines predictions under realistic, constrained sensing conditions.
Comprehensive experiments based on a realistic, field-deployed sensor configuration—augmented with two publicly available benchmark datasets—demonstrate the framework’s effectiveness in leveraging cross-domain data, its robustness to differences in sensor modality, and significant improvement in fatigue detection accuracy under practical, sensor-limited scenarios
1 Introduction
@@ 2 Related Work
2.1 Traditional and Diversified Fatigue Assessment
Traditional fatigue assessment methods are based on subjective evaluation tools, including self-reports and questionnaires. Since the last century, numerous classic fatigue assessment scales have been widely utilized, including the Stanford Sleepiness Scale (SSS) [10], proposed by Stanford University; the Multidimensional Fatigue Inventory (MFI) [11], developed by Smets et al.; and the International Fitness Scale (IFIS) [12], which measures changes in overall vigor and affect. These scales encompass a multitude dimensions of fatigue and serves as the foundation for traditional fatigue assessment. Nevertheless, both questionnaires and self-reports are highly influenced by an individual’s psychological state and external environment, resulting in subjective and inconsistent outcomes [13]. Moreover, most of these methods [14, 15, 16] are post-hoc assessments, making them unsuitable for real-time fatigue monitoring and preventive measures-both of which are essential in high-stakes environments, such as driving [17] and aviation [18].
To address these limitations, researchers have explored more objective measures of human fatigue. In fact, a person’s fatigue state can be reflected in changes in physiological signals, such as heart rate (HR), heart rate variability (HRV), electrocardiogram (ECG), electroencephalogram (EEG), galvanic skin response (GSR), skin yemperature (ST), blood pressure(BP), electromyography(EMG) and blood oxygen saturation (SpO2). Studies have shown that when drivers enter a fatigued state, their HRV metrics change significantly. Notably, a decrease is often observed in the time domain on measures such as how much the time between heartbeats naturally varies and how much consecutive heartbeats differ from each other. This suggests that the ability of the body to regulate heart activity is weakened. Similarly, HRV indicators in the frequency domain, such as the low-frequency and high-frequency ratios, often become unstable, reflecting an imbalance in the regulatory function of the autonomic nervous system [19]. Additionally, GSR signals during fatigue are usually characterized by a rise in baseline conductance level and response fluctuations, which is closely correlated to enhanced sympathetic nerve activities [20]. Through real-time monitoring and joint analysis of these specific indicators, the subtle changes in the driver’s physiological state can be captured more precisely, thus significantly improving the sensitivity and accuracy of the fatigue detection system [21]. With technological advancement, physiological signal acquisition devices are becoming smaller and more user-friendly, significantly driving research in real-time fatigue monitoring [22, 23, 24].
@@ 2.1 Traditional and Diversified Fatigue Assessment
In response to these challenging, researchers have explored non-invasive alternatives, leading to the rise of vision-based fatigue detection.
These methods assess fatigue by analyzing facial features [30], such as eyelid closure time [31], yawning frequency [32], facial feature [33], and head movement [34] from images captured by cameras in real time. Despite their many advantages, these methods also face challenges, such as lighting variations and camera angles, which can affect image quality, and ultimately detection accuracy. In addition, long-term monitoring human subjects’ facial and body movements in work environments raises privacy concerns [35] and may lead to employee resistance. Along with this trend, alternative visual signals, like pupil dilation and gaze tracking, have gained attention in the field of fatigue detection. Eyes are often considered as the most intuitive window of mental state, which can directly reflect the immediate response to fatigue and show high accuracy in fatigue assessment. Studies have shown that pupil diameter and gaze focus are important indicators for assessing workload and fatigue status [36, 37, 38], including our own study [39] to collect human eye and physiological data for combined fatigue detection. However, while pupil dilation and gaze tracking offer promising insights, they are also sensitive to external environmental factors. Pupil dilation can be influenced by ambient lighting conditions, which may reduce its reliability in dynamic or uncontrolled environments. Similarly, gaze tracking can be affected by head movements, occlusions, and variations in camera positioning, limiting its stability and accuracy unless the camera is confined to a controlled environment, such as the one in our study [39] using Hololens.
@@
2.2 Data Synthesis Methods
@@
Early developments in DS primarily leveraged statistical regression techniques, where synthetic data was generated through parameterized modeling of observed data distributions. For instance, Linear regression and Bayesian regression were among the first approaches adopted to model temporal dependencies and fill in missing physiological signals by assuming probabilistic relationships among features [40]. As datasets grew in complexity and dimensionality, regularized models such as ridge regression [41] and lasso regression [42] were introduced to improve generalization. However, despite these advancements, statistical models remained fundamentally limited in their ability to capture long-term dependencies and nonlinear temporal dynamics, which are prevalent in complex physiological signals.
Deep learning, on the other hand, has provided promising solutions to this challenge. Recurrent Neural Networks (RNNs), particularly Long Short-Term Memory (LSTM) networks, have proven effective in capturing long-term temporal dependencies in multivariate time-series data. Several studies have demonstrated their success in a variety of applications, such as GRU-D, which incorporates decay mechanisms and masking for clinical time-series data [43]; an LSTM-based imputation model for ECG signals that outperforms classical methods [44]; and BRITS, which enables bidirectional gradient-based imputation without strong distributional assumptions [45]. In parallel, Convolutional Neural Networks (CNN) have shown great success in extracting local spatial and temporal features. Their strength has inspired hybrid models that combine CNNs with LSTMs to jointly leverage spatial and temporal dependencies for more accurate imputation and signal reconstruction [46]. Additionally, Generative Adversarial Networks (GANs) have emerged as a leading approach for producing high-fidelity synthetic signals that are statistically indistinguishable from real ones. Examples can be found in [47, 48].
@@
2.3 Open Fatigue Detection Datasets
A review of relevant literature has identified several open datasets for fatigue detection. As shown in Table 1, the use of diverse physiological signals reaffirms our observation that researchers tend to select those sensors that are readily accessible and convenient for their studies. Meanwhile, we also note that certain signals are commonly used across diverse datasets, including HR, PPG, EDA, Acceleration(ACC) and ST.
The commonality and diversity across datasets present both challenges and opportunities for fatigue detection. For example, while FatigueSet [49], MEFAR [50], and VPFD [39] share similar types of data, they use different sensors as detailed in a later session - Empatica E4 for the first two datasets and Google Pixle Watch 2 for the third. These variations in data collection methods, sensor specifications, and pre-processing techniques can hinder standardization and generalization. FatigueSet [49] and MEFAR [50], collected under controlled laboratory conditions, include high-fidelity signals—such as EEG in both datasets and ECG in FatigueSet—that are often impractical for real-world deployment. In contrast, the VPFD [39] dataset was designed for real-world use and thus relies on more accessible sensors except for EEG and ECG. This distinction directly motivates our evaluation of the central hypothesis posed in the Introduction: Can fatigue-relevant knowledge learned from high-resolution EEG and ECG sources be effectively transferred to enhance fatigue prediction in settings like VPFD, where only context-appropriate sensors are available? In the following sections, we present the results of this exploration.
Table 1: Summary of fatigue detection datasets and acquisition devices
Dataset Human Hours Data Types
Subjects PPG GSR HR ST ACC EYE EEG ECG EMG BP FACE
CogBeacon [51] 19 35 $\checkmark$ $\checkmark$
CLAS [52] 62 31 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
MePhy [53] 60 8 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
OperEYEVP [54] 10 10 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
WESAD [55] 15 13 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
FatigueSet [49] 12 13 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
MEFAR [50] 23 28 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
VPFD [39] 4 9 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
3 Heterogeneous and Multi-source Fatigue Detection Framework
3.1 Problem Definition
@@
Since the obtained sensors tend to be heterogeous across different domains in the data collection process. Thus, we have
  • 1.
    $\mathcal{X}^{\cap}:=\mathcal{X}_{T}\cap\mathcal{X}_{s}$  (common sensors between $\mathcal{D}_{T}$ and $\mathcal{D}_{s}$ )
  • 2.
    $\mathcal{X}^{s-}:=\mathcal{X}_{s}\setminus\mathcal{X}_{T}^{\cap}$  (sensors present in $\mathcal{D}_{T}$ but missing in $\mathcal{D}_{s}$)
  • 3.
    $\mathcal{X}^{s+}:=\mathcal{X}_{T}\setminus\mathcal{X}_{s}^{\cap}$  (sensors present in dataset $\mathcal{D}_{s}$ but missing in $\mathcal{D}_{T}$)
@@ 3.2 Theoretical Analysis
Through Theorem 1 and Theorem 2, it is evident that without additional sensors $a$, if the distribution distance between $\mathbb{Q}_{a}$ and $\mathbb{Q}_{\hat{a}}$, is less than a constant, i.e, $d_{\mathcal{H}\Delta\mathcal{H}}(\mathbb{Q}_{a},\mathbb{Q}_{\hat{a}})\leq\left|\sqrt{\frac{2R^{2}}{n}}\left(\sqrt{I(a,y)}-\sqrt{I(x,y)}\right)-\Delta-\epsilon_{ideal}\right|$, we can obtain a tighter generalization error bound, thereby improving the generalization performance of the model. This result validates that generating advanced sensors can enhance the performance of fatigue detection.
@@
3.3 Learning Objective
Motivated by Theorem 1 and 2, we provide our heterogeneous and multi-source framework. Specifically, given a target domain $\mathcal{D}_{T}$, our learning objective is to detect fatigue with knowledge transferring from multiple source domains (e.g., high-fidelity knowledge extracted from controlled environments) $\{\mathcal{D}_{s}\}$ for $s=1,\ldots,S$. Here, as a naive attempt, we aims to transfer knowledge by utilizing the remaining domains to generate the missing sensors (i.e., $\mathcal{X}^{s+}$) in $\mathcal{D}_{s}$ The learning objective is defined as follows:
@@ 3.3 Learning Objective
Moreover, motivated by Theorem 3 , a batch normalization step [58], referred as BatchNorm, is applied before fusion, to reduce $d_{\mathcal{H}\Delta\mathcal{H}}(\mathbb{P}_{T}^{\cap},\mathbb{P}_{s}^{\cap})$. Additionally, to optimize the objective function in Equation 5, particularly the task-specific loss, we adopt the Jacobian Norm Regularization method [59]-referred to as JacobianNorm later.
3.4 Framework
The overall framework is summarized in Figure 1, with the detailed procedures outlined in Algorithms 1 and 2. Specifically, we propose a proof-of-concept approach that enables cross-domain knowledge integration for improved fatigue detection in heterogeneous sensor settings. Specifically, the method is designed to handle data collected from different sensor configurations with partially overlapping feature spaces. When new data from a different sensor setup becomes available, the framework compares its feature space to that of the target domain. If discrepancies are found—i.e., missing modalities—the method initiates a targeted imputation process to infer the missing signals. The aligned and augmented data is then incorporated into the model to enhance prediction accuracy in the target domain.
@@ 4.1.1 Datasets
4.1.2 Implementation Environment and Data Preprocessing
@@
4.2 Experimental Design and Results
4.2.1 Regression‑Based Imputation Model Selection
Without loss of generality, the first set of experiments validates our imputer selection by assessing its ability to recover a single blocked modality. We compare three regression architectures—MLP, LSTM (Long Short-Term Memory), and CNN1D—on two datasets: FatigueSet (with ECG masked) and MEFAR (with EEG masked). In each case, the missing channel is replaced by the model prediction, and we record the mean squared error. Table 2 presents these results. The MLP consistently achieves the lowest loss, demonstrating its superiority as an imputer. Consequently, we employ the MLP for all subsequent modality imputation tasks.
Table 2: Regression performance of different models
FatigueSet-ECG MEFAR-EEG

### v2

@@ Abstract
Note to Practitioners—The proposed work is motivated by the growing need for practical and reliable fatigue detection in industrial applications where operator fatigue poses serious safety and productivity risks. While high-performing physiological sensors like EEG provide high-fidelity signals, their use in these environments is limited by practical constraints such as cost, setup complexity, and environment sensitivity. Thus, industry deployment often relies on context-appropriate sensors that may not fully capture fatigue-related states. To resolve this conflict, we propose a multi-source heterogeneous fatigue detection framework that enables field-deployed systems to take advantage of high-fidelity data collected from controlled environments, even if those modalities are unavailable during deployment. Our experiments validate the practicality, robustness, and generalization of our approach, offering a scalable solution to effective fatigue monitoring in sensor-constrained industry settings.
Index Terms:
@@
publicationid: pubid:
I Introduction
@@ II Related Work
II-A Traditional and Diversified Fatigue Assessment
Traditional fatigue assessment methods are based on subjective evaluation tools, including self-reports and questionnaires. Since the last century, numerous classic fatigue assessment scales have been widely utilized, including the Stanford Sleepiness Scale (SSS) [10], proposed by Stanford University; the Multidimensional Fatigue Inventory (MFI) [11], developed by Smets et al.; and the International Fitness Scale (IFIS) [12], which measures changes in overall vigor and affect. These scales encompass a multitude dimensions of fatigue and serve as the foundation for traditional fatigue assessment. Nevertheless, both questionnaires and self-reports are highly influenced by an individual’s psychological state and external environment, resulting in subjective and inconsistent outcomes [13]. Moreover, most of these methods [14, 15, 16] are post hoc assessments, making them unsuitable for real-time fatigue monitoring and preventive measures-both of which are essential in high-stakes environments, such as driving [17] and aviation [18].
To address these limitations, researchers have explored more objective measures of human fatigue. In fact, a person’s fatigue state can be reflected in changes in physiological signals, such as heart rate (HR), heart rate variability (HRV), electrocardiogram (ECG), electroencephalogram (EEG), galvanic skin response (GSR), skin temperature (ST), blood pressure(BP), electromyography(EMG) [19, 20] and blood oxygen saturation (SpO2). Studies have shown that when drivers enter a fatigued state, their HRV metrics change significantly. Notably, a decrease is often observed in the time domain on measures such as how much the time between heartbeats naturally varies and how much consecutive heartbeats differ from each other. This suggests that the ability of the body to regulate heart activity is weakened. Similarly, HRV indicators in the frequency domain, such as the low-frequency and high-frequency ratios, often become unstable, reflecting an imbalance in the regulatory function of the autonomic nervous system [21]. Additionally, GSR signals during fatigue are usually characterized by a rise in baseline conductance level and response fluctuations, which is closely correlated to enhanced sympathetic nerve activities [22]. Through real-time monitoring and joint analysis of these specific indicators, the subtle changes in the driver’s physiological state can be captured more precisely, thus significantly improving the sensitivity and accuracy of the fatigue detection system [23]. With technological advancement, physiological signal acquisition devices are becoming smaller and more user-friendly, significantly driving research in real-time fatigue monitoring [24, 25].
@@ II-A Traditional and Diversified Fatigue Assessment
In response to these challenges, researchers have explored noninvasive alternatives, leading to the increase of vision-based fatigue detection [31]. These methods assess fatigue by analyzing facial features, such as eyelid closure time [32], yawning frequency [33], and body movement [34, 35] from images captured by cameras in real time. Despite their many advantages, these methods also face challenges, such as lighting variations and camera angles, which can affect image quality and ultimately detection accuracy. In addition, long-term monitoring human subjects’ facial and body movements in work environments raises privacy concerns [36] and may lead to employee resistance. Along with this trend, alternative visual signals, like pupil dilation and gaze tracking, have gained attention in the field of fatigue detection. Eyes are often considered as the most intuitive window of the mental state, which can directly reflect the immediate response to fatigue and show high accuracy in fatigue assessment. Studies have shown that pupil diameter and gaze focus are important indicators for assessing workload and fatigue status [37, 38, 39], including our own study [40] to collect human eye and physiological data for combined fatigue detection. However, while pupil dilation and gaze tracking offer promising insights, they are also sensitive to external environmental factors. Pupil dilation can be influenced by ambient lighting conditions, which may reduce its reliability in dynamic or uncontrolled environments. Similarly, gaze tracking can be affected by head movements, occlusions, and variations in camera positioning, limiting its stability and accuracy unless the camera is confined to a controlled environment, such as the one in our study [40] using Hololens.
@@
II-B Data Synthesis Methods
@@
Early developments in data synthesis(DS) primarily leveraged statistical regression techniques, where synthetic data was generated through parameterized modeling of observed data distributions. For instance, Linear regression and Bayesian regression were among the first approaches adopted to model temporal dependencies and fill in missing physiological signals by assuming probabilistic relationships among features [41]. As datasets grew in complexity and dimensionality, regularized models such as ridge regression [42] and lasso regression [43] were introduced to improve generalization. However, despite these advancements, statistical models remained fundamentally limited in their ability to capture long-term dependencies and nonlinear temporal dynamics, which are prevalent in complex physiological signals.
Deep learning, on the other hand, has provided promising solutions to this challenge. Recurrent Neural Networks (RNNs), particularly Long Short-Term Memory (LSTM) networks, have proven effective in capturing long-term temporal dependencies in multivariate time-series data. Several studies have demonstrated their success in a variety of applications, such as GRU-D, which incorporates decay mechanisms and masking for clinical time-series data [44]; an LSTM-based imputation model for ECG signals that outperforms classical methods [45]; and BRITS, which enables bidirectional gradient-based imputation without strong distributional assumptions [46]. In parallel, Convolutional Neural Networks (CNN) have shown great success in extracting local spatial and temporal features. Their strength has inspired hybrid models that combine CNNs with LSTMs to jointly leverage spatial and temporal dependencies for more accurate imputation and signal reconstruction [47]. Additionally, Generative Adversarial Networks (GANs) have emerged as a leading approach for producing high-fidelity synthetic signals that are statistically indistinguishable from real ones. Examples can be found in [48, 49, 50].
@@
II-C Open Fatigue Detection Datasets
A review of relevant literature has identified several open datasets for fatigue detection. As shown in Table I, the use of diverse physiological signals reaffirms our observation that researchers tend to select those sensors that are readily accessible and convenient for their studies. Meanwhile, we also note that certain signals are commonly used across diverse datasets, including HR, PPG, EDA, Acceleration(ACC) and ST.
The commonality and diversity across datasets present both challenges and opportunities for fatigue detection. For example, while FatigueSet [51], MEFAR [52], and VPFD [40] share similar types of data, they use different sensors as detailed in a later section - Empatica E4 for the first two datasets and Google Pixel Watch 2 for the third. These variations in data collection methods, sensor specifications, and pre-processing techniques can hinder standardization and generalization.
FatigueSet [51] and MEFAR [52], collected under controlled laboratory conditions, include high-fidelity signals—such as EEG in both datasets and ECG in FatigueSet—that are often impractical for real-world deployment. In contrast, the VPFD [40] dataset was designed for real-world use and thus relies on more accessible sensors except for EEG and ECG. This distinction directly motivates our evaluation of the central hypothesis posed in the Introduction: Can fatigue-relevant knowledge learned from high-resolution EEG and ECG sources be effectively transferred to enhance fatigue prediction in settings like VPFD, where only context-appropriate sensors are available? In the following sections, we present the results of this exploration.
TABLE I: Summary of fatigue detection datasets and acquisition devices
Dataset Human Subjects Hours Data Types
PPG GSR HR ST ACC EYE EEG ECG EMG BP FACE
CogBeacon [53] 19 35
CLAS [54] 62 31
MePhy [55] 60 8
OperEYEVP [56] 10 10
WESAD [57] 15 13
FatigueSet [51] 12 13
MEFAR [52] 23 28
VPFD [40] 4 9
III Heterogeneous and Multi-source Fatigue Detection Framework
III-A Problem Definition
@@
Since the obtained sensors tend to be heterogeneous across different domains in the data collection process. Thus, we have
  • $\mathcal{X}^{\cap}:=\mathcal{X}_{T}\cap\mathcal{X}_{s}$
    (common sensors between $\mathcal{D}_{T}$ and $\mathcal{D}_{s}$ )
  • $\mathcal{X}^{s-}:=\mathcal{X}_{s}\setminus\mathcal{X}_{T}^{\cap}$
    (sensors present in $\mathcal{D}_{T}$ but missing in $\mathcal{D}_{s}$)
  • $\mathcal{X}^{s+}:=\mathcal{X}_{T}\setminus\mathcal{X}_{s}^{\cap}$
    (sensors present in dataset $\mathcal{D}_{s}$ but missing in $\mathcal{D}_{T}$)
@@ III-B Theoretical Analysis
Through Theorem 1 and Theorem 2, it is evident that without additional sensors $a$, if the distribution distance between $\mathbb{Q}_{a}$ and $\mathbb{Q}_{\hat{a}}$, is less than a constant, i.e., $d_{\mathcal{H}\Delta\mathcal{H}}(\mathbb{Q}_{a},\mathbb{Q}_{\hat{a}})\leq\left|\sqrt{\frac{2R^{2}}{n}}\left(\sqrt{I(a,y)}-\sqrt{I(x,y)}\right)-\Delta-\epsilon_{ideal}\right|$, we can obtain a tighter generalization error bound, thereby improving the generalization performance of the model. This result validates that generating advanced sensors can enhance the performance of fatigue detection.
@@
III-C Learning Objective
Motivated by Theorem 1 and 2, we provide our heterogeneous and multi-source framework. Specifically, given a target domain $\mathcal{D}_{T}$, our learning objective is to detect fatigue with knowledge transferring from multiple source domains (e.g., high-fidelity knowledge extracted from controlled environments) $\{\mathcal{D}_{s}\}$ for $s=1,\ldots,S$. Here, as a naive attempt, we aim to transfer knowledge by utilizing the remaining domains to generate the missing sensors (i.e., $\mathcal{X}^{s+}$) in $\mathcal{D}_{s}$ The learning objective is defined as follows:
@@ III-C Learning Objective
Moreover, motivated by Theorem 3 , a batch normalization step [60], referred to as BatchNorm, is applied before fusion, to reduce $d_{\mathcal{H}\Delta\mathcal{H}}(\mathbb{P}_{T}^{\cap},\mathbb{P}_{s}^{\cap})$. Additionally, to optimize the objective function in Equation 5, particularly the task-specific loss, we adopt the Jacobian Norm Regularization method [61]-referred to as JacobianNorm later.
III-D Framework
The overall framework is summarized in Figure 1, with the detailed procedures outlined in Algorithms 1 and 2. Specifically, we propose a proof-of-concept approach that enables cross-domain knowledge integration for improved fatigue detection in heterogeneous sensor settings. Specifically, the method is designed to handle data collected from different sensor configurations with partially overlapping feature spaces. When new data from a different sensor setup becomes available, the framework compares its feature space to that of the target domain. If discrepancies are found—i.e., missing modalities—the method initiates a targeted imputation process to infer the missing signals. The aligned and augmented data are then incorporated into the model to enhance prediction accuracy in the target domain.
@@ IV-A1 Datasets
IV-A2 Implementation Environment and Data Preprocessing
@@
IV-B Experimental Design and Results
IV-B1 Regression‑Based Imputation Model Selection
Without loss of generality, the first set of experiments validates our imputer selection by assessing its ability to recover a single blocked modality. We compare three regression architectures—MLP, LSTM (Long Short-Term Memory), and CNN1D—on two datasets: FatigueSet (with ECG masked) and MEFAR (with EEG masked). In each case, the missing channel is replaced by the model prediction, and we record the mean squared error. Table II presents these results. The MLP consistently achieves the lowest loss, demonstrating its superiority as an imputer. Consequently, we employ the MLP for all subsequent modality imputation tasks.
TABLE II: Regression performance of different models
Model FatigueSet-ECG MEFAR-EEG